OpenStudy (anonymous):

Find an equation of the tangent line to the ellipse: x^2/a^2+ y^2/b^2 = 1 at the point (x0,y0)?

OpenStudy (wisp):

Do you already have your tangent line?

OpenStudy (anonymous):

Yeah The tangent line is supposed to be (x0x)/a^2 + y0y/b^2

OpenStudy (anonymous):

I can't seem to simplify my answer to get that.

OpenStudy (anonymous):

=1

OpenStudy (wisp):

Wasn't that supposed to be the ellipse?

OpenStudy (anonymous):

Nope. The ellipse is x^2/a^2+ y^2/b^2 = 1. It's really similar though. I'm kinda wondering if my textbook messed up or something.

OpenStudy (wisp):

Are you in Calc AB? Or BC?

OpenStudy (anonymous):

Uh... I'm in IB Math HL, so I'm not sure. What I'm doing is right after Pre-Calc though.

OpenStudy (wisp):

Ahhh, I see. I'm in Calc AB, but I think what my teacher told me might help. The slope of a line (like, say, maybe a tangent line) is -A/B. What's really confusing me about this are the variables that are being used. I'm used to working with the actual numbers instead of a and b.

OpenStudy (anonymous):

Yeah, I know what you mean. I breezed through the number questions, but I just got a bit stuck here.

OpenStudy (wisp):

You might also want to take the derivative of the actual ellipse (though, I'm not sure if they teach that in IB Math; I'm thinking they do?). Then you would find out with the derivative of the ellips is equal to the tangent line's slope (-A/B).

OpenStudy (anonymous):

Yeah, I took the derivative then got\[{2b ^{3}x(a-a'x) + 2ya ^{3}(y'b-b'y)}\div(a ^{3}b ^{3})\]

OpenStudy (anonymous):

I'm trying to simplify it by inserting it into y-y1 = m(x-x1), but it isn't goign so well.

OpenStudy (wisp):

The entire equation should be equal to 0 (the one you just wrote out), and wherever you have y', there is also a dy/dx (the thing you're after). Once you get dy/dx by itself, that's the derivative. Then you find where that equals the slope of your ellipse at -A/B, and your tangent line would be (y-y1)=(dy/dx)(x-x1).

OpenStudy (anonymous):

Ah, I hate expanding. I'll see if that'll work, thanks either way!

OpenStudy (wisp):

Oh, and a and b should be arbitrary, so their derivatives should be zero.

OpenStudy (anonymous):

oh right I forgot that. that should help things.

OpenStudy (anonymous):

so treating a and b as constants instead of variables got me a lot father: I ended up with \[y0(y-y0)\div b ^{2}=x0(x0-x)\div a ^{2}\]

OpenStudy (anonymous):

You think there's any way I can get a 1 out of that mess?

OpenStudy (wisp):

You don't need a 1, right? 1 is just what the original ellipse is equal to. the derivative is the slope at a specific point. Where's your dy/dx?

OpenStudy (anonymous):

The y' is \[(-xb^{2})\div (ya^{2})\]

OpenStudy (wisp):

Alright, and since, at the end of the question, you were asked to find he tangent line (the line of the slope) at the point (x0,y0), you would first plug x0 and y0 into your y'. What y' ends up equaling is the m in your equation (y-y1)=m(x-x1). And x1 and y1 are x0 and y0. That would be your tangent line.

OpenStudy (anonymous):

yup. that's how I got y0(y−y0)÷b2=x0(x0−x)÷a2

OpenStudy (anonymous):

no wait I think I got it

OpenStudy (anonymous):

If I use the original equation, x0^2/a^2 = 1- y0^2/b^2 right?

OpenStudy (anonymous):

So if I plug that into the equation I got above... the two y0^2/b^2 cancel and I'm left with 1!

OpenStudy (wisp):

=) Is your questioned answered now?

OpenStudy (anonymous):

yup thanks :D

OpenStudy (wisp):

No problem, glad I could be of help ^-^.

OpenStudy (anonymous):

Wish I could help you out but your question's way beyond my ability. if no one answers here you could try Yahoo Answers. There's some pretty smart people on there sometimes.

OpenStudy (wisp):

Alright, and don't sweat it ;). I'll go ahead and try that now since It's getting close to 9:30 and I've still got to finish my paper.

OpenStudy (anonymous):

Good luck!